OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 15x times the square root of 3 all divided by the square root of 75x squared?

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ \sqrt{ 75 x^2} }\]